a 10% solution of fertilizer is to be mixed with a 50% solution of fertilizer in order to get 160 gallons of 40% solution. How many gallons of the 10% solution and the 50% solution should be mixed?
Answer by mananth(16949)

let 10 % be x
50 % will be 160-x
Mixture 40% 160 liters
..
Concentration equation
..
0.1x+0.5(160-x)= 160*.4
0.1x +80-0.5x = 64
-0.4x = -80+64
-0.4x = -16
divide by -0.4
x= -16/-0.4
x=40 liters 10% solution